OpenStudy (anonymous):

Solve the equation. 4(x + 2) + 8 = −3(x − 8) − 1

OpenStudy (anonymous):

Are we solving for x here?

OpenStudy (anonymous):

Alright so solve the distributive parts of the equation (the parentheses parts)... then match like terms with each other...

jagr2713 (jagr2713):

Distribute 4 over x+2. 4 (x+2) = 4 x+8: 4 x+8+8 = -1-3 (x-8)

jagr2713 (jagr2713):

then ddd the numbers in 4 x+8+8 together. Add like terms. 8+8 = 16 4 x+16 = -1-3 (x-8)

jagr2713 (jagr2713):

then distribute -3 over x-8. -3 (x-8) = 24-3 x: 4 x+16 = 24-3 x-1

jagr2713 (jagr2713):

then add the numbers in -3 x-1+24 together. Add like terms. 24-1 = 23 4 x+16 = 23-3 x

jagr2713 (jagr2713):

then move terms with x to the left hand side. Add 3 x to both sides: 4 x+3 x+16 = (3 x-3 x)+23

jagr2713 (jagr2713):

then we look for two terms that sum to zero. 3 x-3 x = 0: 4 x+3 x+16 = 23 then after we add like terms in 4 x+3 x+16. 4 x+3 x = 7 x: 7 x+16 = 23

jagr2713 (jagr2713):

then Isolate terms with x to the left hand side. Subtract 16 from both sides: 7 x+(16-16) = 23-16 Look for two terms that sum to zero. 16-16 = 0: 7 x = 23-16

jagr2713 (jagr2713):

then finally evaluate 23-16. 23-16 = 7: 7 x = 7 Divide both sides by a constant to simplify the equation.

jagr2713 (jagr2713):

divide both sides of 7 x = 7 by 7: (7 x)/7 = 7/7

jagr2713 (jagr2713):

Any nonzero number divided by itself is one. 7/7 = 1: x = 7/7 so x =1

OpenStudy (anonymous):

i have to write this down

jagr2713 (jagr2713):

i mean if you want to

OpenStudy (anonymous):

ok because this is alot to comprehend, and i can go back and look at it

jagr2713 (jagr2713):

ok just read it and you will understand

OpenStudy (anonymous):

ok thank you so much

jagr2713 (jagr2713):

yw